# Chaotic-Integrable Transition for Disordered Orbital Hatsugai-Kohmoto Model

### v1

@@ 3.2 Spectral Form Factor
While the level spacing statistic provides valuable GOE, the SFF helps offer supporting information. The ramp feature observed in the SFF is attributed to long-range level repulsion [4, 17]. When balanced against the effects that keep the energy finite, this repulsion gives rise to a less rigid eigenvalue structure. Consequently, due to its low rigidity, the SFF exhibits a ramp-shaped increase that linearly grows over time before transitioning into a plateau. Note that $g(t)$ is dominated by the disconnected SFF $g_{d}(t)$ before the dip time and by the connected SFF $g_{c}(t)$ after the dip time. In Ref. [4], a dip-ramp-plateau behavior was observed for the SYK model, particularly in the ramp and plateau regions, with a subtle difference in early-time behavior. For the disordered orbital HK model, the unfolded SFF patterns in integrable and chaotic cases at infinite temperature fall between the Poisson and GOE predictions of RMT, as shown in Fig. 3.
Figure 3: SFF for infinite temperature, averaged over 50 random samples for the disordered orbital HK model with 8 orbitals with half-filling and 0 total spin. The chaotic phase is characterized by $\langle\tilde{r}\rangle=0.53$ with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, and the integrable phase by $\langle\tilde{r}\rangle=0.387$ with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$. The unfolded SFF is displayed in the upper plot, while the connected portion is shown in the lower plot.
In the case where
$\langle\tilde{r}_{n}\rangle=0.53,$ (25)
@@
Figure 4: SFF at infinite temperature for the disordered orbital HK model with 6 and 8 orbitals at half-filling and zero total spin. The upper plot displays the results for the 6-orbital scenario, averaged across 500 random samples. The lower plot illustrates the 8-orbital scenario, averaged across 50 random samples.
@@ 4 OTOCs
where $N_{o}$ is the number of $n_{\alpha,\sigma}$. We calculate $C(t)$ by averaging over random samples $\langle C(t)\rangle_{t_{h},U}$. As shown in the down plot of Fig. 5, in the Poisson distribution regime where $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$, the OTOCs remain constant across different temperatures. However, in the GOE distribution regime, with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, the late time behavior of OTOCs increases with temperatures, as illustrated in the up plot of Fig. 5. This behavior possibly arises from varying exponents during early times. Due to the numerical difficulties in calculating the Lyapunov exponent, the late-time behavior is particularly noteworthy.
Figure 5: Results for 30 random samples with 6 orbitals. The OTOCs versus time on a log scale at different temperatures. The upper plot corresponds to GOE-level statistics with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, while the lower plot represents Poisson-level statistics with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$. For the Poisson case, the OTOCs calculations are totally temperature independent. For the GOE case, the late-time values of OTOCs are temperature dependent. Note that $N_{o}=12$.
Additionally, as depicted in Fig. 6, the system size does not affect the temperature-dependent late-time behavior of OTOCs. However, when we analyze the plateau values of OTOCs in other many-body models, such as the SYK2 model, the disorder-free SYK model [20], and the SYK4 model, we observe that the saturated values are temperature-independent across both integrable and maximally chaotic models, as illustrated in Fig. 7. Furthermore, we also examine the regularization issue through the regularized OTOCs,
@@
as shown in Fig. 7. The normalized $F(t)$, i.e., $F(t)/F(0)$, shows that the saturated values tend to exhibit similar behavior in the SYK2, disorder-free SYK model, and SYK4 models. This result aligns with the $C(t)$ calculation, suggesting that the relationship between the late-time behavior of OTOCs and temperature may not be a reliable indicator of chaotic properties in many-body systems. However, a different behavior happens in the disordered orbital HK model. Therefore, the late-time OTOCs, in general, depends on the regularization.
Figure 6: $C(t)$ calculation with different system size. (a) 2 orbitals with results averaged over 1000 random samples, (b) 4 orbitals with 100 random samples, (c) 6 orbitals with 30 random samples, (d) 8 orbitals with only 1 random sample. The results show that the system size does not affect the temperature-dependent late-time behavior of OTOCs.
Figure 7: OTOCs calculations: $C(t)$ and normalized $F(t)$ for various models. (a)(b) SYK2 model with $\kappa=1$ and $N=18$ over 10 random samples. (c)(d) Disorder-free SYK model with $N=18$. (e)(f) Disordered orbital HK model with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, considering 6 orbitals over 30 random samples. The number of number operators $N_{o}=12$. (g)(h) SYK4 model with $N=18$ over 10 random samples.
@@ 5 Discussion and Conclusion
In this paper, we investigated the adjacent gap ratio (a measure of energy level repulsion) [14], spectral form factor [16], and out-of-time-order correlators in the disordered orbital HK model. The disordered orbital HK model, when Fourier transformed to position space, shows similarities to the SYK2+SYK4 model [10], a well-known system for studying quantum chaos and integrability. Including the four-point interaction term in the adjacent gap ratio calculations reveals an increase in equity (presumably referring to a trend toward more universal chaotic behavior). SFF and OTOCs also corroborate this observation. SYK2 and disorder-free SYK models exhibit temperature independence in late-time OTOC behavior, similar to the SYK4 model. The disordered orbital HK model demonstrates dependence on regularization in its late-time behavior, suggesting that late-time OTOCs are not a reliable indicator of many-body chaos. This implies that using OTOC saturation values alone to distinguish between integrability and chaos may not always work.
We delved into an intriguing area of quantum chaos, particularly regarding the transition between chaotic and integrable behavior within quantum systems. Studying this transition through simpler models like the disordered orbital HK model is promising because it is less complex than models like the more involved SYK2+SYK4 model. In classical mechanics, integrable systems can be solved exactly, with as many conserved quantities as degrees of freedom, leading to regular, predictable motion. Chaotic systems, on the other hand, exhibit sensitive dependence on initial conditions and complex dynamics. Extending this distinction to quantum systems introduces a challenge, as quantum chaos needs a straightforward analog to classical chaos.
The HK model is a lattice-based system where electrons hop between orbitals in a disordered manner [6, 7], and can be modified with random variables to study NFL behavior and chaos. Unlike the SYK models [1], which are all-to-all interaction models (thus more complex to solve and numerically intensive), the disordered HK model’s simplicity comes from local interactions on the momentum space, allowing us to track the onset of quantum chaos and integrability in a more manageable framework. By tuning the disorder or interaction strength in the disordered orbital HK model, we can induce a transition from integrable to chaotic behavior. This transition is essential for understanding quantum chaos in more complex systems, particularly NFLs. NFL behavior often arises in disordered systems, where quasiparticle descriptions break down. Understanding the chaotic regime could illuminate strongly correlated electron systems and unconventional superconductors. NFLs deviate from the conventional Fermi liquid theory by exhibiting non-quasiparticle excitations. The chaotic nature of some NFLs suggests a deeper connection between chaos and the breakdown of conventional many-body theory. By studying the disordered orbital HK model, which might show NFL behavior under certain conditions, and comparing it with more complex models like SYK2+SYK4, we can explore how chaos influences the NFL state and its transitions. By bridging the chaotic-integrable transition using simpler models, our study can provide fresh insights into quantum chaos, offering a more intuitive understanding of how chaotic dynamics emerge even in systems with relatively few bodies. This has far-reaching implications for quantum information and condensed matter physics.
@@ Acknowledgments
We would like to express our gratitude to Masaki Tezuka for his helpful discussion. CTM would thank Nan-Peng Ma for his encouragement. PYC acknowledges support from the National Science and Technology Council of Taiwan under Grants No. NSTC 113-2112-M-007-019. Both PYC and YLL thank the National Center for Theoretical Sciences, Physics Division for its support. CTM acknowledges the Nuclear Physics Quantum Horizons program through the Early Career Award (Grant No. DE-SC0021892).

### v2

@@ 3.2 Spectral Form Factor
While the level spacing statistic provides valuable GOE, the SFF helps offer supporting information. The ramp feature observed in the SFF is attributed to long-range level repulsion [4, 18]. When balanced against the effects that keep the energy finite, this repulsion gives rise to a less rigid eigenvalue structure. Consequently, due to its low rigidity, the SFF exhibits a ramp-shaped increase that linearly grows over time before transitioning into a plateau. Note that $g(t)$ is dominated by the disconnected SFF $g_{d}(t)$ before the dip time and by the connected SFF $g_{c}(t)$ after the dip time. In Ref. [4], a dip-ramp-plateau behavior was observed for the SYK model, particularly in the ramp and plateau regions, with a subtle difference in early-time behavior. For the disordered orbital HK model, the unfolded SFF patterns in integrable and chaotic cases at infinite temperature fall between the Poisson and GOE predictions of RMT, as shown in Fig. 2. In the case where
$\langle\tilde{r}\rangle=0.531,$ (22)
@@
Figure 2: SFF at infinite temperature for the disordered orbital HK model with 6, 8, and 10 orbitals at half-filling and zero total spin. The chaotic phase is characterized by $\langle\tilde{r}\rangle=0.531$ with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, and the integrable phase by $\langle\tilde{r}\rangle=0.386$ with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$. (a) The results for the 6-orbital scenario are averaged across 500 random samples. (b) The 8-orbital scenario, averaged across 50 random samples. (c) The 10-orbital scenario, averaged across 5 random samples. (d) The connected portion of the 10-orbital scenario.
@@ 4 OTOCs
where $N_{o}$ is the number of $n_{\alpha,\sigma}$. We calculate $C(t)$ by averaging over random samples $\langle C(t)\rangle_{t_{h},U}$. As shown in Fig. 3, in the Poisson distribution regime where $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$, the OTOCs remain constant across different temperatures. However, in the GOE distribution regime, with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, the late time behavior of OTOCs increases with temperatures, as illustrated in Fig. 3. This behavior possibly arises from varying exponents during early times. Due to the numerical difficulties in calculating the Lyapunov exponent, we only consider the late-time behavior of OTOC.
Figure 3: Results for 30 random samples with 6 orbitals. The OTOCs versus time on a log scale at different temperatures. (a) GOE-level statistics with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$. (b) Poisson-level statistics with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=0$. For the Poisson case, the OTOCs calculations are totally temperature independent. For the GOE case, the late-time values of OTOCs are temperature dependent. $C(t)$ calculation with different system sizes: (c) 2 orbitals with results averaged over 1000 random samples; (d) 4 orbitals with 100 random samples; (e) 6 orbitals with 30 random samples; (f) 8 orbitals with only 1 random sample. The results show that the system size does not affect the temperature-dependent late-time behavior of OTOCs.
Additionally, as depicted in Fig. 3, the system size does not affect the temperature-dependent late-time behavior of OTOCs. However, when we analyze the plateau values of OTOCs in many many-body models, such as the SYK2 model, the disorder-free SYK model [21], and the SYK4 model, we observe that the saturated values are temperature-independent across both integrable and maximally chaotic models, as illustrated in Fig. 4. Furthermore, we also examine the regularization issue through the regularized OTOCs,
@@
as shown in Fig. 4. The normalized $F(t)$, i.e., $F(t)/F(0)$, shows that the saturated values tend to exhibit similar behavior in the SYK2, disorder-free SYK, and SYK4 models. This result aligns with the $C(t)$ calculation, suggesting that the relationship between the late-time behavior of OTOCs and temperature may not be a reliable indicator of chaotic properties in many-body systems. However, a different behavior happens in the disordered orbital HK model. Therefore, the late-time OTOCs, in general, depend on the regularization.
Figure 4: OTOCs calculations: $C(t)$ and normalized $F(t)$ for various models. (a)(b) Disordered orbital HK model with $\langle t_{h}^{2}\rangle=1$ and $\langle U^{2}\rangle=3.3$, considering 6 orbitals over 30 random samples. The number of number operators $N_{o}=12$. (c)(d) SYK2 model with $\kappa=1$ and $N=18$ over 10 random samples. (e)(f) Disorder-free SYK model with $N=18$. (g)(h) SYK4 model with $N=18$ over 10 random samples.
@@ 5 Discussion and Conclusion
In this paper, we investigated the adjacent gap ratio (a measure of energy level repulsion) [15], spectral form factor [17], and out-of-time-order correlators in the disordered orbital HK model. The disordered orbital HK model, when Fourier transformed to position space, shows similarities to the SYK2+SYK4 model [11], a well-known system for studying quantum chaos and integrability. SYK2 and disorder-free SYK models exhibit temperature independence in late-time OTOC behavior, similar to the SYK4 model. The disordered orbital HK model demonstrates dependence on regularization in its late-time behavior, suggesting that late-time OTOCs are not a reliable indicator of many-body chaos. This implies that using OTOC saturation values alone to distinguish between integrability and chaos may not always work.
We delved into an intriguing area of quantum chaos, particularly regarding the transition between chaotic and integrable behavior within quantum systems. Studying this transition through simpler models like the disordered orbital HK model is promising because it is less complex than models like the more involved SYK2+SYK4 model. In classical mechanics, integrable systems can be solved exactly, with as many conserved quantities as degrees of freedom, leading to regular and predictable motions. Chaotic systems, on the other hand, exhibit sensitive dependence on initial conditions and complex dynamics. Extending this distinction to quantum systems introduces a challenge, as quantum chaos needs a straightforward analog to classical chaos.
The HK model is a lattice-based system where electrons hop between orbitals in a disordered manner [6, 7], and can be modified with random variables to study NFL behavior and chaos. By tuning the disordered interaction strength in the disordered orbital HK model, we can induce a transition from integrable to chaotic behavior. This transition is essential for understanding quantum chaos in more complex systems, particularly NFLs. NFL behavior often arises in strongly correlated systems, where quasiparticle descriptions break down. The chaotic nature of some NFLs suggests a deeper connection between chaos and the breakdown of conventional many-body theory. By studying the disordered orbital HK model, which might show NFL behavior under certain conditions, and comparing it with other chaotic models like SYK2+SYK4, we can explore how chaos influences the NFL state and its transitions.
@@ Acknowledgments
We would like to express our gratitude to Masaki Tezuka for his helpful discussion. CTM would thank Nan-Peng Ma for his encouragement. PYC acknowledges support from the National Science and Technology Council of Taiwan under Grant No. NSTC 113-2112-M-007-019. Both PYC and YLL thank the National Center for Theoretical Sciences, Physics Division for their support. YLL also thanks Chi-Ting Ho for insightful discussions. Computational resources were provided by Academia Sinica Grid Computing Centre, supported by Grant No. AS-CFII-112-103.